OpenStudy (anonymous):

Help me! Soybean meal is 16% protein,cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320lbs mixture that is 12% protein?

OpenStudy (anonymous):

Making the assumption that only soybean meal and cornmeal are added to the mixture. We get x = amount of soybean meal and y = amount of cornmeal x+y = 320. solving for y we get y =320 -x. We also know that 0.16(x) + .08(y) = 320*(.12) or .16x + 0.08(320-x) = 38.4 simplify 0.08x + 25.6 = 38.4 0.08x = 12.8 x = 160 subsituting into x+y = 320 we get y = 160 So lbs of cornmeal = 160 and lbs of soybeam meal is 160